Find f(x + 1) when f(x) = (x^2 -4) / (x+4)
Answer by rothauserc(4718)

f(x) = (x^2 -4) / (x+4)
:
f(x+1) = ( (x+1)^2 - 4 ) / (x+1+4)
:
f(x+1) = ( x^2 + 2x + 1 - 4 ) / (x+5)
:
f(x+1) = ( x^2 +2x -3 ) / (x+5)
:
*********************************
f(x+1) = ((x+3) * (x-1)) / (x+5)
*********************************
: